OpenStudy (maeganr):

How do I solve each system of inequalities by graphing? 3x + y < 1 x - y < 3 Both are supposed to be less than or equal to. The less than signs should have a line under both of them.

OpenStudy (anonymous):

Sketch 3x+y=1 and x-y=3 for me please.

OpenStudy (maeganr):

idk

OpenStudy (anonymous):

Then how are you supposed to do these types of questions when you don't know how to sketch lines? Aren't you meant to learn the basics before continuing?

OpenStudy (anonymous):

If you don't know how to sketch lines, why didn't you ask your teacher for help when he/she was teaching you how to sketch lines?

OpenStudy (maeganr):

yeah. but my teacher sucks.

OpenStudy (anonymous):

Your teacher does not "suck" if you're brave enough to actually go up to the teacher and ask for help....

OpenStudy (maeganr):

i do.

OpenStudy (anonymous):

Then why are you saying you can't sketch lines right now?

OpenStudy (maeganr):

because she hasnt even taught us that.

OpenStudy (anonymous):

If she hasn't taught you how to sketch lines such as this: y=x+1 then you must be kidding yourself that you should be asking this question on this website because to be straight forward with you, you're not going to be able to answer this question unless your teacher teaches you how to sketch lines such as what I put up as an example.

OpenStudy (maeganr):

Thank you. I appreciate your help. ...

OpenStudy (anonymous):

I didn't help you with anything but to advise you that your teacher must teach you the basics of sketching before giving you questions like this. I hope you take that on board with anything you do because, if you don't have the basics, how are you supposed to excel at maths?

OpenStudy (maeganr):

ikr.